# What Language Models Know But Don’t Say: Non-Generative Prior Extraction for Generalization

### v1

@@ A.3.1 Feature Synonyms
We evaluated the impact of feature synonyms on model performance to determine whether changing feature names could introduce bias based on word selection. To do this, we created three sets of synonyms for the features in our evaluation datasets. Table 7 presents examples of these features and their corresponding synonyms.
Dataset Original Feature Synonym Set 1 Synonym Set 2
Blood recency Time since last donation Days elapsed since previous donation
frequency Number of donations made Donation count
Heart Age Patient age Years of life
Cholesterol Serum cholesterol level Blood cholesterol concentration
MaxHR Peak heart rate Highest heart rate reached
ExerciseAngina Angina during exercise Exercise-provoked chest pain
Diabetes Glucose Blood glucose level Plasma glucose concentration
BMI Body mass index Weight-to-height ratio
DiabetesPedigree Family diabetes history score Genetic diabetes risk
Table 7: Representative examples of feature name variations tested in synonym ablation study (Synonym Set 0 = original feature names).
Results in Table 8 reveal minimal performance variation across three distinct feature naming schemes, with AUC ranges of 0.007, 0.002, and 0.001 for blood, heart, and diabetes datasets respectively. These negligible differences demonstrate that our elicitation method is robust to superficial linguistic variations in feature descriptions. This stability is significant because it indicates the LLM is capturing genuine semantic understanding of the features’ predictive relationships rather than relying on brittle pattern matching or memorized associations with specific phrasings.
Dataset Syn0 Syn1 Syn2 Range
Bank 0.692 0.692 0.696 0.004
Blood 0.657 0.654 0.655 0.004
Credit Score 0.755 0.755 0.755 0.000
Diabetes 0.836 0.835 0.835 0.000
Heart 0.887 0.887 0.887 0.000
Income 0.844 0.842 0.844 0.002
Liver 0.667 0.666 0.667 0.001
Occupancy 0.988 0.988 0.988 0.001
Jungle 0.79 0.778 0.778 0.012
Pima 0.836 0.836 0.837 0.000
Mean Range 0.0013
Table 8: Synonym Ablation Study: Robustness to feature name variations across all datasets. Three synonym sets are tested.
A.3.2 Feature Description
We analyzed the impact of adding further description to the features and the target could help the LLM arrive at better estimates of the prior distributions.
Table 9 reveals minimal and inconsistent impact from including feature names alongside descriptions in prompts, with a mean absolute AUC difference of only 0.0084 (0.84%) across all 10 datasets. Performance changes are bidirectional, with 5 datasets showing marginal improvements (range: +0.002 to +0.038) and 5 showing marginal declines (range: -0.003 to -0.016) when feature descriptions are included. This 50/50 split and small standard deviation (0.0135) indicate no systematic benefit or detriment to either prompt format, suggesting the LLM extracts equivalent semantic information regardless of whether features are referenced by name-description pairs or names alone. The method’s robustness to this formatting choice demonstrates that its effectiveness stems from semantic understanding rather than sensitivity to linguistic structure.
Dataset w/o Desc w/ Desc $\Delta$
Bank 0.692 0.673 -0.019
Blood 0.652 0.650 -0.002
Credit Score 0.756 0.756 +0.000
Diabetes 0.836 0.836 +0.000
Heart 0.886 0.893 +0.007
Income 0.845 0.853 +0.008
Liver 0.668 0.663 -0.005
Occupancy 0.989 0.986 -0.002
Jungle 0.778 0.777 -0.000
Pima 0.837 0.835 -0.002
Mean $|\Delta|$ 0.0050
Table 9: Feature Description Ablation Study: Impact of including feature descriptions in prompts across all datasets. Comparing name-only prompts vs. name+description format.
@@ A.4 Datasets and OOD Splits
A.4.1 Datasets
We evaluate our prior elicitation method on ten binary, and multi-class classification datasets:
@@ A.4.1 Datasets
    Bank Marketing Moro et al. (2014): Predicts whether a client will subscribe to a term deposit based on direct marketing campaign data from a Portuguese banking institution. Features include client demographics (age, job, marital status, education), campaign information (contact type, number of contacts, previous campaign outcome), and economic indicators (employment variation rate, consumer price index). The binary classification task identifies clients likely to subscribe to term deposits.
@@
    Blood Transfusion Yeh (2008b): Predicts whether a blood donor will donate within a given time window using data from the Blood Transfusion Service Center in Hsin-Chu City, Taiwan. Features are based on the RFMTC marketing model: Recency (months since last donation), Frequency (total number of donations), Monetary (total blood donated in c.c.), and Time (months since first donation). The binary classification task identifies donors likely to donate in March 2007.
@@
    Give Me Some Credit Kaggle (2011): Predicts the probability of financial distress in the next two years using credit and demographic data from Kaggle. Features include checking account status, credit duration, credit history, credit amount, savings account balance, employment status, and age. The binary classification task evaluates financial risk for 1,000 borrowers.
  4. 4.
    Diabetes Islam et al. (2020): Predicts the likelihood of diabetes based on self-reported symptoms and demographic information. Features include age, gender, polyuria, polydipsia, sudden weight loss, weakness, polyphagia, genital thrush, visual blurring, itching, irritability, delayed healing, partial paresis, muscle stiffness, alopecia, and obesity. The dataset contains 520 instances collected via questionnaires and is formulated as a binary classification task.
  5. 5.
    Heart Disease Janosi et al. (1988): Predicts heart disease using clinical and demographic attributes such as age, sex, resting blood pressure, cholesterol level, fasting blood sugar, maximum heart rate, exercise-induced angina, ST depression (oldpeak), chest pain type, resting ECG results, and ST slope. The binary classification task identifies individuals with high likelihood of heart disease.
  6. 6.
    Adult Income Becker and Kohavi (1996a): Predicts whether an individual’s annual income exceeds \$50K based on 1994 U.S. Census data. Features include age, education level, marital status, occupation, relationship status, workclass, sex, native country, capital gain/loss, and hours worked per week. The binary classification task uses demographic and employment data from 48,842 records.
  7. 7.
    Chess Endgame Shapiro (1983): Predicts game outcomes in King-Rook vs. King-Pawn chess endgame positions. Features encode piece positions using strength rankings, file coordinates, and rank coordinates for white and black pieces. The binary classification task determines winning positions in this chess endgame with 44,820 board configurations.
  8. 8.
    Indian Liver Patient Ramana and Venkateswarlu (2012): Predicts whether a patient has liver disease using clinical measurements collected from northeast Andhra Pradesh, India. Features include age, gender, total bilirubin, direct bilirubin, alkaline phosphatase, alanine aminotransferase (SGPT), aspartate aminotransferase (SGOT), total proteins, albumin, and albumin/globulin ratio. The binary classification task distinguishes liver patients from healthy individuals using data from 583 records (416 liver patients, 167 non-liver patients).
  9. 9.
    Occupancy Detection Candanedo (2016a): Predicts room occupancy status from environmental sensor measurements in an office setting. Features include temperature (Celsius), relative humidity (%), light (Lux), CO2 concentration (ppm), humidity ratio (kg water-vapor/kg-air), and timestamp information. The binary classification task detects occupancy using measurements recorded at one-minute intervals.
  10. 10.
    Pima Indians Diabetes Smith et al. (1988): Predicts the onset of diabetes within five years in females of Pima Indian heritage over age 21. Features include number of pregnancies, glucose level, blood pressure, skin thickness, insulin level, BMI, diabetes pedigree function, and age. The binary classification task uses medical measurements from 768 women from a population near Phoenix, Arizona with high diabetes incidence rates.
Dataset Method Mean Mag. Std (Uncertainty) Std Ratio Performance
[…]
AutoElicit 0.051 0.058 $1\times$
[…]
AutoElicit 0.015 0.124 $1\times$
[…]
AutoElicit 0.017 1.379 $1\times$
[…]
AutoElicit 0.008 0.047 $1\times$
[…]
AutoElicit 0.015 0.087 $1\times$
[…]
AutoElicit 0.070 0.044 $1\times$
[…]
AutoElicit 0.085 0.059 $1\times$
[…]
AutoElicit 0.007 0.122 $1\times$
[…]
AutoElicit 0.025 0.114 $1\times$
[…]
AutoElicit 0.041 0.024 $1\times$
Table 10: Comparison of LoID and AutoElicit across datasets in terms of magnitude, uncertainty, and performance.
A.4.2 OOD Splits
We construct out-of-distribution (OOD) train splits to simulate a training set that does not represent the entire population. For each dataset, we consider numerical features with at least two unique values and apply five shift strategies to induce distributional gaps between training and test data:
  1. 1.
    extreme_10, training on the bottom 10%
  2. 2.
    extreme_5_95, training on the bottom 5%
  3. 3.
    moderate_20_80, training on the bottom 20%;
  4. 4.
    tail_0_50, training on the 0–25% range;
  5. 5.
    tail_50_100, training on the 50–75% range.
For each split, we designate a different feature as the shift variable (cycling through available features), compute the corresponding quantile thresholds, and construct train–test masks while enforcing a minimum of 50 samples per split. Models are trained on the OOD training set and evaluated on the entire dataset to quantify performance degradation under distribution shift. Splits with insufficient class diversity (fewer than two classes in either the training or test set) are discarded. This procedure simulates realistic deployment settings in which models must generalize to regions of the feature space not observed during training.
For Heart Disease, we use the cholesterol split (moderate_20_80), training on the 20–80% quantile range and testing on extreme values, resulting in an AUC drop from 0.933 to 0.870 ($\Delta=0.063$). Liver Disease uses the alkaline phosphatase split (tail_0_50), training on the bottom 50% and testing on the top 50%, with degradation from 0.768 to 0.698 ($\Delta=0.070$). Blood Donation employs the recency split (extreme_10_90), training on donors in the bottom 10% and testing on the top 10%, experiencing the largest degradation from 0.755 to 0.579 ($\Delta=0.176$). Occupancy Detection uses the light intensity split (moderate_20_80), although this split exhibits insufficient class diversity for full evaluation. Diabetes and Pima both use the pregnancies split (extreme_10_90), training on the bottom 10% and testing on the top 10%, with degradation from 0.839 to 0.813 ($\Delta=0.027$). CreditG partitions on credit amount (extreme_5_95), training on the bottom 5% and testing on the top 5%, yielding a drop from 0.786 to 0.710 ($\Delta=0.076$). Income uses the education level split (extreme_5_95), with the steepest decline from 0.908 to 0.781 ($\Delta=0.127$). Bank Marketing splits on call duration (extreme_10_90), experiencing severe degradation from 0.888 to 0.508 ($\Delta=0.380$), representing the most extreme covariate shift in our evaluation. Finally, Jungle Chess uses the black piece strength split (tail_0_50), training on the 0–50% range and testing on 50–100%, with degradation from 0.810 to 0.712 ($\Delta=0.098$). These splits simulate realistic deployment scenarios in which models trained on one subpopulation must generalize to substantially different data distributions.
[…]
Bank 0.51 0.61 0.72 0.61 0.89 +30.0
Blood 0.58 0.63 0.67 0.62 0.75 +31.3
Credit Score 0.71 0.74 0.76 0.76 0.79 +0.0
Diabetes 0.81 0.82 0.84 0.81 0.84 +104.5
Heart 0.87 0.91 0.90 0.92 0.93 -31.8
Income 0.78 0.81 0.85 0.83 0.91 +14.9
Liver 0.70 0.68 0.69 0.70 0.77 -9.9
Occupancy 0.98 0.72 1.00 0.75 1.00 +1155.7
Jungle 0.71 0.72 0.79 0.75 0.81 +40.0
Pima 0.82 0.81 0.84 0.81 0.84 +98.1
Average 0.74 0.75 0.80 0.76 0.86 +143.3
Table 11: Comparison of LoID priors and priors with 0 means and stds in the Goldilock zone on OOD datasets. $\Delta$ Gap Closed shows the percentage point difference in gap closed between LoID and Prior = $\mathcal{N}=(0,0.45)$. OOD-LR is logistic regression trained on OOD data (lower bound), Cap is logistic regression trained on full IID data (upper bound).
A.5 LoID vs Existing Baselines
A.5.1 Prior Performance
Since LoID and AutoElicit employ similar prior elicitation strategies and exhibit more comparable performance than other baselines, Table 10 focuses on analyzing where and why each method performs better. The underlying intuition behind these differences is discussed in detail in Section 5.1.
A.5.2 Computational Efficiency

### v2

@@ A.3.1 Feature synonyms
We evaluated the impact of feature synonyms on model performance to determine whether changing feature names could introduce bias based on word selection. To do this, we created three sets of synonyms for the features in our evaluation datasets. Table 5 presents examples of these features and their corresponding synonyms. We test three variants: (1) concise, (2) verbose, and (3) clinical feature descriptions. Examples of these variants are provided in Table 5.
Results
Table 6 shows the effect of different grammatical styles for describing features. We find that Feature description style shows modest and inconsistent effects. Across four datasets, we observe an average performance range of 5.7 AUC points between the best and worst description variants. Variant 3 (clinical/formal) achieves the highest mean performance (0.707), outperforming concise (0.679) and verbose (0.664) styles. However, the effect is dataset-dependent: Adult Income shows substantial sensitivity (13.1 point range), with Variant 2 performing poorly (0.636 vs. 0.753/0.767 for other variants), while Stroke exhibits minimal sensitivity (1.9 point range). For medical datasets (Stroke, Heart Failure), clinical terminology appears marginally beneficial, suggesting domain-appropriate phrasing may improve prior extraction. The inconsistency across datasets indicates that while LLMs are not entirely robust to paraphrasing, the sensitivity is unpredictable and relatively small.
Variant Example Prompt
Variant 1 Consider a person who is older. Is this person at risk of stroke? Answer:
Consider a person who has hypertension. Is this person at risk of stroke? Answer:
Consider a person who has fewer months since last donation. Does this person have blood donation likelihood? Answer:
Variant 2 Consider a person who is of older age. Is this person at risk of stroke? Answer:
Consider a person who has a diagnosis of hypertension. Is this person at risk of stroke? Answer:
Consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
Variant 3 Consider a person who exhibits advanced age. Is this person at risk of stroke? Answer:
Consider a person who presents with hypertension. Is this person at risk of stroke? Answer:
Consider a person who demonstrates recent donation activity. Does this person have blood donation likelihood? Answer:
Table 5: Example prompts for feature description variants. All variants use the same template structure but differ in how features are described.
A.3.2 Contextual preambles
We test whether providing domain-knowledge context before feature-effect questions improves prior extraction. For each feature, we prepend an explanatory preamble describing the causal mechanism (e.g., “Stroke risk increases significantly with age due to arterial stiffening and accumulated vascular damage. Based on this medical knowledge,”) before asking the standard question. Table 7 shows example prompts with and without contextual preambles.
Table 6 shows the effect of adding domain-knowledge context before feature-effect questions.
Feature Description Variants Contextual Preambles
Dataset Variant 1 Variant 2 Variant 3 Range $\checkmark$ $\times$ $\Delta$
(Concise) (Verbose) (Clinical) (Baseline)
Stroke 0.685 0.703 0.704 0.019 0.694 0.707 +0.014
Blood Donation 0.567 0.582 0.593 0.026 0.583 0.530 -0.053
Heart Failure 0.713 0.736 0.764 0.051 0.752 0.638 -0.114
Adult Income 0.753 0.636 0.767 0.131 0.745 0.771 +0.026
Mean 0.679 0.664 0.707 0.057 0.693 0.662 -0.032
Table 6: Ablation study results. Left: Effect of feature description style. Right: Effect of contextual preambles.
Condition Example Prompt
No Preamble (Baseline) Consider a person who is older. Is this person at risk of stroke? Answer:
Consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
Consider a person who has higher ejection fraction. Does this person have poor outcome? Answer:
With Preamble Stroke risk increases significantly with age due to arterial stiffening and accumulated vascular damage. Based on this medical knowledge, consider a person who is older. Is this person at risk of stroke? Answer:
People who have donated blood recently are typically more engaged and committed donors. Based on this behavioral pattern, consider a person who has donated blood more recently. Does this person have blood donation likelihood? Answer:
Higher ejection fraction indicates better cardiac pump function and improved prognosis. Based on this cardiac physiology, consider a person who has higher ejection fraction. Does this person have poor outcome? Answer:
Table 7: Example prompts comparing baseline (no preamble) with contextual preamble condition.
Results
Contextual preambles degrade performance on average. Adding domain-knowledge explanations before feature-effect questions reduces mean AUC by 3.2 points (0.693 → 0.662). The effect is highly inconsistent: Adult Income (+2.6 points) and Stroke (+1.4 points) benefit modestly from preambles, while Heart Failure (-11.4 points) and Blood Donation (-5.3 points) suffer substantial degradation. This suggests that explicit reasoning chains interfere with the LLM’s implicit token probability distributions rather than enhancing them. Preambles may bias the model toward generating coherent text completions (which favor the preamble’s stated relationship) rather than revealing the underlying probability mass over answer tokens. For prior extraction via logit probabilities, minimal prompting appears superior to verbose contextual framing.
@@ A.4 Datasets and OOD splits
A.4.1 Public datasets
Adult Income Becker and Kohavi (1996): Predicts whether an individual’s annual income exceeds \$50K using demographic and educational features such as age, work class, education, marital status, occupation, race, sex, and hours worked per week. The binary classification task is based on 48,842 records from the U.S. Census Bureau Adult dataset.
Cervical Cancer Fernandes et al. (2017): Predicts the risk of cervical cancer using demographic factors, sexual health history, and medical test results. Features include age, number of sexual partners, age at first intercourse, smoking habits, hormonal contraceptive use, and HPV test results. The dataset contains 858 patient records collected from clinics in Colombia.
Chronic Kidney Disease Rubini et al. (2015): Classifies the presence of chronic kidney disease using clinical and laboratory measurements, including blood pressure, glucose level, blood urea, serum creatinine, hemoglobin, and albumin. The binary classification task is based on 400 patient records from multiple hospitals.
Bank Marketing Moro et al. (2014): Predicts whether a client will subscribe to a term deposit based on demographic attributes and call-related features such as age, job type, marital status, education, balance, contact communication type, and duration of the last call. The dataset contains 45,211 records collected from Portuguese banking campaigns.
Blood Transfusion Service Center Yeh (2008): Predicts whether a donor will donate blood again using historical donation behavior, including recency, frequency, monetary amount of previous donations, and time since first donation. The dataset consists of 748 donor records from a blood transfusion center in Taiwan.
Stroke Fedesoriano (2020): Predicts the risk of stroke using demographic and health-related features such as age, hypertension, heart disease, glucose level, BMI, smoking status, and gender. The binary classification task is based on 5110 patient records collected from hospitals in India.
Heart Failure UCI Machine Learning Repository (2020): Predicts mortality due to heart failure using clinical measurements including age, ejection fraction, serum creatinine, serum sodium, platelets, and smoking status. The dataset contains 299 patient records collected from hospitals in Portugal.
Diabetes Prediction Mustafatz (2023): Classifies diabetes risk using blood glucose levels, BMI, and other demographic and health-related features. The dataset contains 520 patient records and is used to predict the onset of diabetes in individuals.
Give Me Some Credit Kaggle (2011): Predicts the probability of financial distress within two years using demographic and credit-related features such as age, monthly income, number of past delinquencies, number of open credit lines, and debt-to-income ratio. The dataset consists of 150,000 records from U.S. financial institutions.
Airfoil Self-Noise Brooks et al. (1989): Predicts the noise levels generated by an airfoil using measurements collected by NASA. Features include frequency, angle of attack, chord length, free-stream velocity, and suction side displacement thickness. The regression task is based on 1,503 measurements of airfoil noise under different operating conditions.
Bike Sharing Fanaee-T (2013): Predicts hourly bicycle rental demand based on temporal and weather-related features, including hour of the day, day of the week, temperature, humidity, wind speed, and season. The dataset consists of 17,379 hourly records collected from a bike-sharing system in Washington, D.C.
Combined Cycle Power Plant Tfekci and Kaya (2014): Predicts net hourly electrical energy output using ambient variables such as temperature, pressure, humidity, and exhaust vacuum. The regression task contains 9,568 records collected from a combined cycle power plant over six years.
Concrete Compressive Strength Yeh (1998): Predicts the compressive strength of concrete using mixture components such as cement, blast furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate, as well as curing age. The dataset contains 1,030 records of concrete mixtures tested under controlled conditions.
Communities and Crime Redmond (2002): Predicts per-capita violent crime rates in U.S. communities using demographic, socioeconomic, and law enforcement features, including population, income, employment, education, and police presence. The dataset contains 1,994 community records compiled from U.S. government sources.
A.4.2 Private datasets
Head and Neck Cancer Recurrence Dataset
In collaboration with Henry Ford Health, we compiled a binary classification dataset for Head and Neck Cancer Recurrence. This dataset includes comprehensive features about patient demographics, tumor characteristics, treatment regimens, and clinical staging, and predicts the likelihood of cancer recurrence following radiotherapy. The dataset encompasses 508 patients with head and neck cancers, capturing clinical variables including comorbidity indices (Charlson score), smoking history (pack-years), HPV status, primary tumor site localization, and detailed radiation therapy parameters such as treatment duration, prescription dose, and the critical time interval between biopsy and treatment initiation.
This dataset presents a significant class imbalance (390 non-recurrence vs. 118 recurrence cases), reflecting real-world clinical distributions. The dataset is particularly valuable for evaluating model performance under distributional shift, as we created an out-of-distribution test set focusing on elderly patients ($age\geq 75$), representing a clinically meaningful subpopulation that is not present in any LLM pre-training corpora.
Moreover, this dataset highlights why our approach is necessary. Due to its limited sample size, some patterns learned from the data are not medically plausible. Table 8 shows several noticeable relationships between features and the target variable. For instance, the data suggests that marijuana use is associated with a lower risk of Head and Neck Cancer recurrence, which contradicts established medical evidence Kim et al. (2020). This discrepancy is likely due to sampling bias or confounding factors in the small dataset.
LoID priors address this issue by incorporating domain knowledge into the model. In this case, they can enforce a prior that reflects a more realistic relationship between marijuana use and cancer recurrence. As a result, the model is guided away from spurious correlations, making it more reliable and better suited for generalization to real-world populations.
Recurrence Yes No
Sample Size 118 390
CONTINUOUS FEATURES (Mean [Standard Deviation])
Pack Years 41.3 (38.1) 27.3 (28.7)
Biopsy-RT Duration 44.6 (28.2) 46.2 (47.5)
Average Income 82348.7 (32140.4) 90597.7 (36881.5)
CATEGORICAL FEATURES (Mode [%])
Clinical N 2a (50.0) 0 (29.0)
Clinical T 2 (53.7) 2 (57.6)
Radiation Type IMRT Standard (50.4) IMRT Vmat (52.2)
Alcohol Frequent (41.3) No Alcohol Use (45.8)
Smoking History Former Use (45.4) Never Smoker (41.5)
Grade of Differentiation Moderate (68.7) Moderate (79.0)
Neck RT Bilateral (85.1) Bilateral (91.7)
Tobacco Yes (54.5) No (63.2)
Marijuana No (87.6) No (73.5)
First Primary Site = Larynx-supraglottis No (73.6) No (82.0)
First Primary Site = Larynx-glottis No (82.6) No (92.0)
First Primary Site = Oropharynx-base of tongue No (78.5) No (69.5)
First Primary Site = Oropharynx-tonsil No (79.3) No (70.5)
Concurrent Chemotherapy No (90.1) No (95.5)
Table 8: Notable feature–target relationships in the Head and Neck Cancer Recurrence dataset. While some associations, such as those involving tobacco use, pack years, and alcohol consumption, align with clinical expectations, others are misleading. In particular, a larger proportion of marijuana users appear in the non-recurrence group compared to the recurrence group, which may incorrectly suggest a protective effect. A standard logistic regression model may learn this spurious correlation, despite it lacking medical validity.
A.5 Out-of-Distribution splits
Our goal in this research is to evaluate how well a model trained on a small, specific population generalizes to a larger population. To simulate this, we create conditions where only a subset of the full population is available for training, but the aim remains to extract meaningful patterns from the data. Specifically, we first generate train and test sets representing the entire population using a standard 80-20 shuffle split. From the training set, we then select a subset that excludes certain samples entirely, creating a subpopulation that is not fully representative of the overall population. This setup reflects real-world scenarios, such as Head and Neck Cancer Recurrence, where the limited number of training samples can hinder pattern discovery. Table 9 summarizes the split criteria for all datasets.
Binary Classification:
  • Adult (Income Prediction): Training set excludes individuals age <28<28.
  • Chronic Kidney Disease: Training set excludes patients with blood urea <61.75<61.75 mg/dL.
  • Cervical Cancer: Training set excludes women whose age at first sexual intercourse $\leq 15$.
  • Blood Donation: Training set excludes donors with recency >14>14 months (i.e., recent donors are excluded).
  • Stroke: Training set includes only patients age $\leq 32$.
  • Heart Failure: Training set includes only patients with ejection fraction $\geq 30\%$.
  • Diabetes Prediction: Training set includes patients with blood glucose $\leq 126$ mg/dL; testing set includes hyperglycemic patients with glucose $\geq 126$ mg/dL.
  • Bank Marketing: Training set excludes calls with duration >127>127 seconds.
  • Head & Neck Cancer (HNC): Training set includes patients age $\leq 75$.
  • Good Credit Score: Training set excludes individuals age >31>31.
Regression:
  • Airfoil Noise: Training set includes low-frequency conditions (frequency <1600<1600 Hz).
  • Bike Sharing: Training set includes early hours (hour $\leq 6$).
  • Combined Cycle Power: Training set includes moderate temperatures (AT $\leq 13.51^{\circ}$C).
  • Communities & Crime: Training set includes racially diverse areas (racePctWhite $\leq 0.98$).
  • Concrete Strength: Training set includes high-cement mixes (>350>350 kg/m³), where mixture interactions change.
Dataset Type Split Feature OOD Test Condition $P_{\text{IID}}$ $P_{\text{OOD}}$
Binary Classification
Adult Binary Age Age << 28 0.809 0.414
Cervical Cancer Binary First sexual intercourse Age << 15 0.870 0.440
Chronic Kidney Disease Binary Blood urea BU << 61.75 mg/dL 0.993 0.822
Diabetes Prediction Binary Blood glucose Glucose >> 126 mg/dL 0.963 0.883
Stroke Binary Age Age >> 32 0.842 0.608
Heart Failure Binary Ejection fraction EF << 30% 0.849 0.746
Bank Marketing Binary Call duration Duration >> 127 sec 0.878 0.780
Blood Donation Binary Recency Recency >> 14 months 0.748 0.572
Head & Neck Cancer Binary Age Age >> 75 0.998 0.542
Good Credit Score Binary Age Age >> 31 0.872 0.800
Regression
Concrete Strength Regression Cement Cement >> 350 kg/m³ 98.1 469.2
Communities Crime Regression Race (White %) White >> 98% 0.007 0.171
Airfoil Noise Regression Frequency Freq << 1600 Hz 23.7 77.2
Combined Cycle Power Regression Temperature Temp >> 13.51°C 21.2 962.9
Bike Sharing Regression Hour Hour >> 6 19030 77858
Table 9: Out-of-distribution split specifications for all datasets. Train/test splits are created by partitioning on a single feature to induce distributional shift with non-overlapping support.
A.5.1 Rationale
These splits target realistic distribution shifts where:
@@ A.5.1 Rationale
    Subpopulation exclusion is plausible: Training data may naturally lack certain demographics (young adults in income surveys), extreme operating conditions (high-frequency turbine tests), or rare clinical presentations (severe heart failure cases).
@@
    Feature relationships may differ: Age-income correlations differ for young adults vs. middle-aged workers; thermodynamic equations behave differently at temperature extremes.
@@
    Prior knowledge should help: LLMs may have generalizable knowledge (e.g., "older age increases stroke risk") that transfers to the OOD regime, even if training data provides no direct evidence.
A.6 LoID vs. correlation analysis
In addition to the baselines mentioned in Table 1, we test whether simple correlation analysis might capture similar relationships to LLM priors. To address this, we test a correlation-based baseline that uses feature-target Pearson correlations from the training data as prior means (with fixed $\sigma=1$). Table 10 compares LoID against correlation-based priors. As expected, correlations learned from OOD train sets can mislead the model. In the Bank Marketing dataset where LoID failed because it is dataset-specific and in Cumbine Cycle Power where $\mathcal{N}(0,1)$ could also regain a lot of the performance showing that the train set was actually kinda representing the test set and so correlation studies also got a good AUC. These results further confirm our previous findings.
Binary Classification (AUC) Regression (MSE)
Dataset LoID Corr. $\Delta$ Dataset LoID Corr. $\Delta$
Adult 0.725 0.529 -0.196 Airfoil Noise 60.5 51.4 -9.1
Bank Marketing 0.652 0.770 +0.118 Bike Sharing 78924 69969 -8955
Blood Donation 0.663 0.574 -0.089 Combined Cycle Power 961.9 24.1 -937.8
Cervical Cancer 0.681 0.450 -0.231 Communities & Crime 0.07 0.18 +0.11
Chronic Kidney Disease 0.901 0.575 -0.326 Concrete Strength 121.0 2292.2 +2171.2
Diabetes Prediction 0.906 0.873 -0.033
Heart Failure 0.753 0.742 -0.011
HNC 0.696 0.566 -0.130
Stroke 0.692 0.554 -0.138
Good Credit Score 0.760 0.376 -0.382
Mean 0.741 0.626 -0.142 Mean 16013 14467 -1546
Table 10: LoID (LLM priors) vs. correlation-based priors across 13 datasets (8 classification, 5 regression). For regression, lower MSE is better; negative $\Delta$ indicates correlation outperforms LoID.
A.7 Computational efficiency